Calculate the pressure experienced by a driver at a depth of 40m of sea water. Density of sea water is 1.3x10^3


1
Expert's answer
2021-03-17T18:31:32-0400
"P=\rho gh=1.3\cdot10^3\ \dfrac{kg}{m^3}\cdot9.8\ \dfrac{m}{s^2}\cdot40\ m=5.1\cdot10^5\ Pa."